Shift report 20030205 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Feb 05 2003 - 07:54:29 EST

  • Next message: Flemming Videbaek: "new phobos paper on spectra"
    ------------------------------------------------------------
    Date: Wed Feb  5 07:54:28 2003, Shift: 00:00-08:00
    Supervisor: ZY, with: 
    Log book#13 pages 148 - 148
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam on from 0:15 to 3:40. 
    2. Data runs (MRS at 45 degree A1050, FS at 8 degree 1/8B)
       Run 7790 (trig. 3 212K, trig. 6 51K)
       Run 7795 (trig. 3 239K, trig. 6 57K)
       
       Pedestal runs:
       Run 7796--TPC pedestal run
       Run 7797--non-TPC pedestal run
       Run 7798--DVM run
    3. No detector problem.
    4. For MRS 45 deg. A1050, need another 50K
       For FS 8 deg. 1/8B, another 22K needed according to the plan.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Feb 05 2003 - 07:55:24 EST